# Towards Solving the Gilbert-Pollak Conjecture via Large Language Models

### v1

@@ 6 Conclusion
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
Appendix A Mathematical Preliminaries
A.1 Background on the Gilbert-Pollak Conjecture
A.1.1 Basic Properties of Steiner Minimal Tree
@@
Lemma 11.
Let $S$ be a Steiner Minimal Tree for a set $V$. The following properties hold: 1. Every Steiner point in $S$ has a degree of exactly 33, and the three incident edges meet at $120^{\circ}$ angles. 2. If $|V|=n$, then $S$ contains at most n2n-2 Steiner points. A tree with exactly n2n-2 Steiner points is called a Full Steiner Tree (FST). 3. Any SMT can be decomposed into a union of edge-disjoint Full Steiner Trees. Consequently, it suffices to consider only Full Steiner Trees when proving lower bounds on the Steiner ratio.
@@
Let $V$ be the set of regular points, and let $S$ be $\mathrm{SMT}(V)$. To complete the induction step, we focus on a terminal substructure of $S$. Consider a leaf $A$ of maximum depth in $S$. Let $X$ be the parent of $A$. Since $A$ is at maximum depth, the sibling of $A$ must also be a terminal, denoted as $B$. Let $Y$ be the parent of $X$. The sibling component $D$ incident to $Y$ cannot have a height exceeding that of $X$; consequently, $D$ is necessarily either a single terminal or a Steiner point connected to two terminals.
@@
A.1.2 Core Lemmas
In this section, We will provide a more detailed description of how induction works.
To complete the induction step, we utilize a reduction argument introduced by Du & Hwang (1983).
@@ A.1.2 Core Lemmas
Let $S=\mathrm{SMT}(V)$ be the Steiner Minimal Tree for a set $V$ of $n$ points. Consider a partition of the edges of $S$ into two disjoint sets, $S^{-}$ and $S_{rem}$ (where $S_{rem}=S\setminus S^{-}$). Let $V^{*}\subsetneq V$ be a non-empty subset of regular points such that no point in $V^{*}$ is incident to $S_{rem}$. We define two auxiliary graphs: $t^{*}$: A forest interconnecting $V^{*}$ and $V\setminus V^{*}$ using only vertices in $V$. It must satisfy: (1) For every $x\in V^{*}$, there exists a path in $t^{*}$ connecting $x$ to some $y\in V\setminus V^{*}$. (2) There are no paths in $t^{*}$ connecting any two distinct points $y_{1},y_{2}\in V\setminus V^{*}$. $S^{+}$: A graph (may contain auxiliary vertices not in $V$) such that $S^{+}\cup S_{rem}$ forms a valid Steiner tree (not necessarily minimal) for $V\setminus V^{*}$. Inductive Step: Suppose that for all point sets with fewer than $n$ points, the Steiner ratio is at least $\rho^{\prime}$. If there exists a tuple $(V^{*},S^{-},S^{+},t^{*})$ satisfying $(L_{S^{-}}-L_{S^{+}})/L_{t^{*}}\geq\rho^{\prime}$, then the Steiner ratio for $V$ also satisfies $L_{S}(V)/L_{m}(V)\geq\rho^{\prime}$.
Intuitively, the constraint on $t^{*}$ guarantees that any spanning tree of $V\!\setminus\!V^{*}$ combined with $t^{*}$ yields a spanning tree of $V$.
Based on Lemma 12, we reformulate the verification of a candidate lower bound $\rho$ for the Steiner ratio as a minimax condition. Let $\mathcal{W}$ denote the space of all edge-length vectors associated with the tree topology shown in Figure 10 (${\bm{w}}\in\mathcal{W}$ is something like $(b,c,d,s,\cdots)$).
@@
For any choice of splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$ and geometric configuration ${\bm{w}}\in\mathcal{W}$, we define the splitting function $F_{\tau}({\bm{w}},\rho)$ as: $F_{\tau}({\bm{w}},\rho)=\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}.\vskip-5.0pt$ Let $\mathcal{F}$ be a family of functions corresponding to admissible splits $\tau$. The lower bound $\rho_{\text{Steiner}}\geq\rho$ is established if: $\textstyle\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0.$ (3) We say that a value $\rho$ is feasible if the condition equation 3 holds.
@@
A.2 Geometric Properties of 4-Point Steiner Trees
@@
By construction, points $A$, $U$, $B$ are arranged in counterclockwise order and form an equilateral triangle, as are points $C$, $V$, $D$. Consequently, $\triangle ABU$ and $\triangle CVD$ are equilateral triangles.
@@
Appendix B Our Theoretical Contributions
@@
B.1 The Type A and B Predicate Framework
@@
Lemma 8 (restated). For any $0\leq u<s\leq v<\min(6,n)$, let $H$ be the foot of the perpendicular from $A_{u}$ to line $A_{v}A_{v+1}$. If $H$ lies on the ray $A_{v}A_{v+1}$ and $|A_{u}H|<a_{s}$, then $A_{n}$ will be trapped inside polygon $A_{u}A_{u+1}\ldots A_{v}HA_{u}$.
@@
Consider a simple polygonal chain $X_{1}X_{2}\dots X_{n}$ in the plane, formed by $n$ segments connected end-to-end, satisfying the following conditions: 1. The angle between any two consecutive segments $X_{i}X_{i+1}$ and $X_{i+1}X_{i+2}$ is $120^{\circ}$. 2. When traversing the chain in the order $X_{1},X_{2},\dots,X_{n}$, every turn at each vertex $X_{i}$ is a right turn. If $n\geq 7$ and the ray starting from $X_{6}$ through $X_{7}$ intersects one of the segments $X_{1}X_{2}$, $X_{2}X_{3}$, $\dots$, $X_{4}X_{5}$, then all subsequent vertices $X_{7},X_{8},\dots,X_{n}$ lie within the closed region $\Omega$ enclosed by the segments $X_{1}X_{2},X_{2}X_{3},X_{3}X_{4},X_{4}X_{5},X_{5}X_{6}$ and the line passing through $X_{6}$ and $X_{1}$.
@@
B.2 Proof of Theorems in Section 4
@@
A series of $F\in\mathcal{F}_{\text{ver}}$ can be derived by establishing one of the following two types of lemmas: 1. Trapped Regular Point Lemma: A proposition asserting that when the parameters $\bm{w}$} satisfy certain linear constraints, an implicit regular point is guaranteed to lie within a specific bounded polygonal region. 2. Valid 4-Point Steiner Tree Lemma: A proposition asserting that when $\bm{w}$} satisfies certain orthogonally convex conditions, a specific 4-point Steiner tree structure is valid.
@@
Consequently, for the first type of lemma, let the trapped regular point be denoted by $v_{0}$. We can consider the set $t^{*}$ that contains $v_{0}$ and its corresponding splitting function:
@@
Lemma  15 (restated).The length of the (AB)(AB)-$(CD)$ type Steiner tree is precisely given by $|UV|$, where $U$ denote the position of point $B$ after segment $AB$ is rotated counterclockwise $60^{\circ}$ about point $A$ and $V$ denote the position of point $D$ after segment $CD$ is rotated counterclockwise $60^{\circ}$ about point $C$.
Proof.
By construction, points $A$, $U$, $B$ are arranged in counterclockwise order and form an equilateral triangle, as are points $C$, $V$, $D$. Consequently, $\triangle ABU$ and $\triangle CVD$ are equilateral triangles.
$|AB|\cdot|US|=|AU|\cdot|BS|+|BU|\cdot|AS|.$
$|US|=|AS|+|BS|.$
$|VT|=|CT|+|DT|.$
@@
    Theorem 6 (Vertex-Maximization Property) is the foundation of our entire verification strategy. It fundamentally addresses the challenge of computational tractability by reducing the verification problem over a continuous parameter domain to a check on its discrete vertices [^1] The algorithm implementing this vertex-checking strategy is provided in Appendix D.1.
@@
    • Trapped Regular Point Lemma: Theorem 9 formally establishes this pathway. It leverages the A/B predicates from Section B.1 to transform the search for such lemmas into a formal logical deduction task.
@@
B.3 Proof of Convexity for Splitting Functions
@@ B.3.2 Convexity of $L_{t^{*}}$
Case 1: $|V^{\prime}|\leq 2$. If $|V^{\prime}|=2$, $L_{S^{+}}$ is simply the Euclidean distance between the two points. If $|V^{\prime}|<2$, the length is 0. In both cases, the convexity follows the same logic as Section B.3.2.
@@ B.3.4 Conclusion
B.4 Theoretical Foundation for the Monotonicity Check
In the Feasibility Verification Algorithm, we encounter unbounded hyperrectangles where some edge length parameters extend to $\infty$. To handle these regions without evaluating limits explicitly, we verify that the splitting function $F({\bm{w}},\rho)$ is monotonically non-increasing with respect to the unbounded parameter. This section establishes the geometric foundations and the algorithmic implementation of this check.
@@ B.4.1 Geometric Foundations
B.4.2 Special Analysis on the Monotonicity of $f$
@@
Based on Lemma 21, when verifying monotonicity with respect to variable $f$, we can apply the monotonicity check procedure to the splits generated by the LLM using the 4-Point Steiner Tree Lemmas. This forms a key foundation for the reduction strategy for unbounded $f$ in Appendix E.2.
The vertex-checking strategy of Theorem 6 is not directly applicable to unbounded parameter regions. To address this case, we must first verify the function’s monotonicity along any unbounded dimensions, thereby reducing the problem to one on a bounded domain. The lemmas presented in this section (Lemmas 19, 20, and 21) provide the necessary theoretical guarantees for this critical monotonicity check.
B.5 Correctness Proof for Algorithm 2
$\frac{\partial F}{\partial x}=\rho\frac{\partial L_{t^{*}}}{\partial x}+\frac{\partial L_{S^{+}}}{\partial x}-\frac{\partial L_{S^{-}}}{\partial x}.$
We analyze the bounds using the variables computed in Algorithm 2:
  • By Lemma 19, $\frac{\partial L_{t^{*}}}{\partial x}\leq N_{t^{*}}$.
  • By Lemma 20, $\frac{\partial L_{S^{+}}}{\partial x}\leq\delta_{S^{+}}$.
Substituting these into the derivative:
$\textstyle\frac{\partial F}{\partial x}$ $\leq\rho\cdot N_{t^{*}}+\delta_{S^{+}}-\delta_{S^{-}}\leq N_{t^{*}}+\delta_{S^{+}}-\delta_{S^{-}}.$ (5)
The algorithm returns True only if $N_{t^{*}}+\delta_{S^{+}}\leq 1$ and $\delta_{S^{-}}=1$ (or if all terms are 0).

### v2

@@ 5 Conclusion
Two design principles in our system, we believe, transfer to other mathematical-discovery problems.
Construct a structured reasoning space. Rather than tasking the LLM with end-to-end proofs, we decompose the search into a compositional space of parameterized lemmas. The LLM proposes and simplifies candidates; symbolic tools verify each one. This division of labor plays to the LLM’s strength in creative search while delegating formal correctness to tools that cannot hallucinate.
Reflect on localized verifier failures. We summarize each verifier failure into a structured reflection signal—a minimal sub-region of the search space where the current set of lemmas is insufficient—and feed it back to the LLM. This search–verify–reflect loop should generalize to any setting that admits an exact verifier and whose failures can be localized to a bounded sub-problem.
6 Limitations and Future Work
We highlight two limitations of the current framework, each pointing to a concrete direction for future work.
Symbolic verification does not scale gracefully. Our verifier relies on Cylindrical Algebraic Decomposition (CAD), whose cost grows rapidly with the number of variables and the algebraic degree of the constraints. As a consequence, we adopt a conservative policy in which high-complexity 4-point lemmas are excluded from regimes with unbounded parameters, since CAD over variable parameters becomes intractable in those cases. Extending the inductive structure to more edges and points—or moving to 5-point or larger Steiner-tree topologies—would amplify this bottleneck. Lighter-weight certificates (e.g., interval arithmetic, sum-of-squares relaxations, or learned proof sketches that invoke CAD only on hard sub-regions) are a natural next step.
Human verification remains in the loop. Although the final lower bound is mathematically certified by symbolic tools, every LLM-proposed lemma is manually checked for correctness before being instantiated as a verification function. The framework is therefore not fully autonomous. Reducing this dependence—via formal proof assistants such as Lean, or via stronger automated lemma checkers—is required for end-to-end autonomous mathematical discovery.
Acknowledgements
LW is supported by National Science Foundation of China (NSFC92470123, NSFC62276005) and the State Key Laboratory of General Artificial Intelligence.
@@ Acknowledgements
This paper advances the use of machine learning for open mathematical research. The cost of our experiments is modest—on the order of a few hundred dollars per discovery—suggesting that LLM-assisted theoretical research is accessible to researchers without large compute budgets. We note two broader considerations. First, as autonomous reasoning agents are scaled to many problems in parallel, their aggregate compute and energy footprint becomes non-trivial and deserves explicit accounting. Second, the paradigm of LLMs proposing structured artifacts that are then checked by exact verifiers is dual-use across any domain with a strong verifier (formal theorem proving, software verification, program synthesis); we encourage future work to remain transparent about which steps are mechanically verified and which still rely on human judgment.
@@
Appendix A Related Works
Solving Math Problems with LLMs. Recent advancements in LLMs have demonstrated remarkable potential in mathematical reasoning, extending from standard benchmarks to competition-level challenges. In specialized domains, Trinh et al. (2024) introduced AlphaGeometry, a neuro-symbolic system capable of solving IMO geometry problems without human demonstrations. To tackle broader formal reasoning, HyperTree Proof Search (Lample et al., 2022) pioneered the integration of Transformer-based policies with Monte-Carlo Tree Search (MCTS). Building on this direction, recent works leverage Reinforcement Learning (RL) within formal environments: AlphaProof (Google DeepMind, 2024) achieves silver-medal standards in Lean, while DeepSeek-Prover (Xin et al., 2024, 2025) employs RL to establish state-of-the-art performance in open-ended theorem proving. To enhance reasoning reliability, InternLM-Math (Ying et al., 2024) leverages process supervision to verify intermediate steps. On the infrastructure front, Yang et al. (2023) developed LeanDojo, an open environment enabling large-scale training of these systems. Most recently, Goedel-Prover (Lin et al., 2025, 2026) was introduced as a frontier model, achieving top performance among open-weight models on the MiniF2F benchmark.
AI for Scientific Discovery. Artificial Intelligence has catalyzed paradigm shifts across various scientific disciplines. In biology, AlphaFold (Jumper et al., 2021; Senior et al., 2020; Krokidis et al., 2025) revolutionized structural biology by predicting protein structures with atomic accuracy. In materials science, Merchant et al. (2023) leveraged graph neural networks to discover 2.2 million new crystals (GNoME), dramatically expanding the scale of stable materials known to humanity. Similarly, GraphCast (Lam et al., 2023) demonstrated that AI can model complex physical dynamics to outperform traditional weather forecasting. Moving towards algorithmic discovery, AlphaTensor (Fawzi et al., 2022) utilized reinforcement learning to discover matrix multiplication algorithms surpassing human-designed heuristics. More recently, the focus has shifted towards fully automated research frameworks; for instance, Lu et al. (2024); Yamada et al. (2025) proposed “The AI Scientist”, where LLM-based agents independently generate ideas, execute experiments, and write manuscripts. Following this trajectory, we extend the exploratory capabilities of AI to the domain of Theoretical Mathematics, specifically targeting combinatorial optimization problems.
Appendix B Mathematical Preliminaries
B.1 Background on the Gilbert-Pollak Conjecture
B.1.1 Basic Properties of Steiner Minimal Tree
@@
Lemma 11 (Gilbert & Pollak (1968)).
Let $S$ be a Steiner Minimal Tree for a set $V$. The following properties hold: 1. Every Steiner point in $S$ has a degree of exactly $3$, and the three incident edges meet at $120^{\circ}$ angles. 2. If $|V|=n$, then $S$ contains at most n2n-2 Steiner points. A tree with exactly n2n-2 Steiner points is called a Full Steiner Tree (FST). 3. Any SMT can be decomposed into a union of edge-disjoint Full Steiner Trees. Consequently, it suffices to consider only Full Steiner Trees when proving lower bounds on the Steiner ratio.
@@
Let $V$ be the set of regular points, and let $S$ be $\mathrm{SMT}(V)$. To complete the induction step, we focus on a terminal substructure of $S$. Consider a leaf $A$ of maximum depth (choosing an arbitrary point as root) in $S$. Let $X$ be the parent of $A$. Since $A$ is at maximum depth, the sibling of $A$ must also be a terminal, denoted as $B$. Let $Y$ be the parent of $X$. The sibling component $D$ incident to $Y$ cannot have a height exceeding that of $X$; consequently, $D$ is necessarily either a single terminal or a Steiner point connected to two terminals.
@@
Definition 12.
Given a local structure in Figure 10, let $n$ be the number of edges excluding the normalized edge $AX$. The parameter space $\mathcal{W}$ is defined as $0,+\infty)^{n}$, where the $i$-th entry represents the length of the $i$-th edge. (Therefore ${\bm{w}}\in\mathcal{W}$ is something like $(b,c,d,s,\cdots)$.)
Since we only consider full Steiner trees, for any ${\bm{w}}\in\mathcal{W}$, the positions of all points in the local structure are uniquely determined and depend affinely on $\bm{w}$}.
B.1.2 Core Lemmas
In this section, we provide a detailed formulation of the induction argument underlying Theorem 4. We first formalize the notion of a splitting, which was introduced informally in Section 2.2. We then state the reduction argument of Du & Hwang (1983) in terms of this notion.
Definition 13.
Given a Steiner tree $S$ for a set of points $V$, a tuple $\tau=(V^{*},S^{-},S^{+},t^{*})$ is called a splitting if:
  1. 1.
    $V^{*}\subsetneq V$ is a non-empty subset of $V$.
  2. 2.
    $S^{-}$ is a subset of the edges of $S$. Let $S_{rem}=S\setminus S^{-}$. No point in $V^{*}$ is incident to $S_{rem}$.
  3. 3.
    $S^{+}$ is a graph (may contain auxiliary vertices not in $S$) such that $S^{+}\cup S_{rem}$ forms a valid Steiner tree for $V\setminus V^{*}$.
  4. 4.
    $t^{*}$ is a forest interconnecting $V^{*}$ and $V\setminus V^{*}$ using only vertices in $V$. It must satisfy: (a) For every $x\in V^{*}$, there exists a path in $t^{*}$ connecting $x$ to some $y\in V\setminus V^{*}$. (b) There are no paths in $t^{*}$ connecting any two distinct points $y_{1},y_{2}\in V\setminus V^{*}$.
@@ B.1.2 Core Lemmas
Let $S=\mathrm{SMT}(V)$ be the Steiner Minimal Tree for a set $V$ of $n$ points. Suppose that for all point sets with fewer than $n$ points, the Steiner ratio is at least $\rho^{\prime}$. If there exists a splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$ satisfying $(L_{S^{-}}-L_{S^{+}})/L_{t^{*}}\geq\rho^{\prime}$, then the Steiner ratio for $V$ also satisfies $L_{S}(V)/L_{m}(V)\geq\rho^{\prime}$.
Based on Lemma 14, we reformulate the verification of a candidate lower bound $\rho$ for the Steiner ratio as a minimax condition.
@@
For any splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$ and geometric configuration ${\bm{w}}\in\mathcal{W}$, we define the splitting function $F_{\tau}({\bm{w}},\rho)$ as: $F_{\tau}({\bm{w}},\rho)=\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}.\vskip-5.0pt$ Let $\mathcal{F}=\{F_{\tau}:\tau\text{ is a splitting}\}$ denote the collection of all splitting functions. The lower bound $\rho_{\text{Steiner}}\geq\rho$ is established if: $\textstyle\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0.$ (3) We say that a value $\rho$ is feasible if the condition equation 3 holds.
@@
B.2 Geometric Properties of 4-Point Steiner Trees
@@
By construction, points $A$, $U$, $B$ are arranged in clockwise order and form an equilateral triangle, as are points $C$, $V$, $D$. Consequently, $\triangle ABU$ and $\triangle CVD$ are equilateral triangles.
@@
Appendix C Our Theoretical Contributions
@@
C.1 The Type A and B Predicate Framework
@@
Type A and B Predicate Framework Lemma 8 (restated). For any $0\leq u<s\leq v<\min(6,n)$, let $H$ be the foot of the perpendicular from $A_{u}$ to line $A_{v}A_{v+1}$. If $H$ lies on the ray $A_{v}A_{v+1}$ and $|A_{u}H|<a_{s}$, then $A_{n}$ will be trapped inside polygon $A_{u}A_{u+1}\ldots A_{v}HA_{u}$.
@@
Consider a simple polygonal chain $X_{1}X_{2}\dots X_{n}$ in the plane, formed by $n$ segments connected end-to-end, satisfying the following conditions:
  1. 1.
    The angle between any two consecutive segments $X_{i}X_{i+1}$ and $X_{i+1}X_{i+2}$ is $120^{\circ}$.
  2. 2.
    When traversing the chain in the order $X_{1},X_{2},\dots,X_{n}$, every turn at each vertex $X_{i}$ is a right turn.
If $n\geq 7$ and the ray starting from $X_{6}$ through $X_{7}$ intersects one of the segments $X_{1}X_{2}$, $X_{2}X_{3}$, $\dots$, $X_{4}X_{5}$, then all subsequent vertices $X_{7},X_{8},\dots,X_{n}$ lie within the closed region $\Omega$ enclosed by the segments $X_{1}X_{2},X_{2}X_{3},X_{3}X_{4},X_{4}X_{5},X_{5}X_{6}$ and the line passing through $X_{6}$ and $X_{1}$.
@@
C.2 Proof of Theorems in Section 3
@@
A series of $F\in\mathcal{F}_{\text{ver}}$ can be derived by establishing one of the following two types of lemmas: 1. Trapped Regular Point Lemma: A proposition asserting that when the parameters $\bm{w}$} satisfy certain linear constraints, an implicit regular point is guaranteed to lie within a specific bounded polygonal region. Formally, let $v_{0}$ denote the trapped regular point in the lemma. Then for each splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$, we can derive a verification function $\tilde{F}_{\tau}$ from the splitting function $F_{\tau}$, as long as $t^{*}$ contains $v_{0}$ as an endpoint of an edge. 2. Valid 4-Point Steiner Tree Lemma: A proposition asserting that when $\bm{w}$} satisfies certain orthogonally convex conditions, a specific 4-point Steiner tree structure is valid. Formally, let $X_{1},X_{2},X_{3},X_{4}$ denote the 4 points in the lemma. Then for each splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$, we can derive a verification function $\tilde{F}_{\tau}$ from the splitting function $F_{\tau}$, as long as $S^{+}=\text{SMT}(X_{1},X_{2},X_{3},X_{4})$.
@@
Consequently, for the first type of lemma, let the trapped regular point be denoted by $v_{0}$. We can consider the set $t^{*}$ that contains $v_{0}$ as an endpoint of an edge and its corresponding splitting function:
@@
    Theorem 6 (Vertex-Maximization Property) is the foundation of our entire verification strategy. It fundamentally addresses the challenge of computational tractability by reducing the verification problem over a continuous parameter domain to a check on its discrete vertices [^1] The algorithm implementing this vertex-checking strategy is provided in Appendix F.1..
@@
    • Trapped Regular Point Lemma: Theorem 9 formally establishes this pathway. It leverages the A/B predicates from Section C.1 to transform the search for such lemmas into a formal logical deduction task.
@@
C.3 Proof of Convexity for Splitting Functions
@@ C.3.2 Convexity of $L_{t^{*}}$
Case 1: $|V^{\prime}|\leq 2$. If $|V^{\prime}|=2$, $L_{S^{+}}$ is simply the Euclidean distance between the two points. If $|V^{\prime}|<2$, the length is 0. In both cases, the convexity follows the same logic as Section C.3.2.
@@ C.3.4 Conclusion
C.4 Theoretical Foundation for the Monotonicity Check
The vertex-checking strategy of Theorem 6 is not directly applicable to unbounded parameter regions. To address this case, we must first verify the function’s monotonicity along any unbounded dimensions, thereby reducing the problem to one on a bounded domain. The lemmas presented in this section (Lemmas 21, 22, and 23) provide the necessary theoretical guarantees for this critical monotonicity check.
@@ C.4.1 Geometric Foundations
C.4.2 Special Analysis on the Monotonicity of $f$
@@
Based on Lemma 23, when verifying monotonicity with respect to variable $f$, we can apply the monotonicity check procedure to the splits generated by the LLM using the 4-Point Steiner Tree Lemmas. This forms a key foundation for the reduction strategy for unbounded $f$ in Appendix D.2.
Appendix D Experimental Setup
D.1 Baseline Verification Functions
Figure 11: Sample of Subtree: $D$ is regular
Du & Hwang (1983) Lemma 2
Lemma 24 (Du & Hwang (1983) Lemma 2).
Corollary 25.
Corollary 26.
If $f\leq d$, there must exist an implicit regular point $F$ such that its distance to the terminal is bounded. The bound depends on whether $D$ is a regular point or a Steiner point: If $D$ is a regular point, $|DF|\leq\max\{|DZ|,|DQ|\}$. If $D$ is a Steiner point, $|VF|\leq\max\{|VY|,|VZ|,|VQ|\}$.